OpenStudy (anonymous):

factor the binomial as a difference of two squares. -169p^2 + 4/169

OpenStudy (anonymous):

\[\frac{4}{169}=\left(\frac{2}{13}\right)^2\]

OpenStudy (anonymous):

(2/13)^2-(13p)^2

OpenStudy (anonymous):

and \(-169p^2=-(13p)^2\)